# Fisher-Geometric Diffusion in Stochastic Gradient Descent: Optimal Rates, Oracle Complexity, and Information-Theoretic Limits

### 02 Mar 2026 v1

@@ 8 Numerical Validation
The numerical experiments are small and diagnostic by design. Each is chosen so that (i) a clean theoretical benchmark is available (Lyapunov/OU or discrete Lyapunov), and (ii) the benchmark isolates exactly the mechanism emphasized by the theory: Fisher/Godambe-aligned diffusion, anisotropy, and propagation under curvature. Rather than reporting large tables of means and standard errors, we let the figures carry the evidence; the data-generating scripts record all replicate-level outputs for reproducibility.
Reproducibility.
Table 2 summarizes the experimental parameters. All experiments use $d=2$, Euler–Maruyama discretization, and $R=500$ independent replicates per configuration. Code is available in the supplementary material.
@@
3 (Rotation smoking gun) 2 $\phi\in[0,\pi/2]$ (theory)
@@
The experiments are organized around three claims. Each claim is followed by experiments that validate it and a “so what” connecting it to an OR design implication.
@@ 8.1 Claim 1: The Lyapunov Equilibrium Law Is Quantitatively Predictive
where $H\succ 0$ is the local curvature, $F\succeq 0$ is the Fisher/Godambe noise shape, $\eta$ is a stepsize scale, and $b$ is the batch size. We choose a two-dimensional sharp/flat curvature $H=\mathrm{diag}(1,0.1)$ and take $F=I_{2}$, then sweep $b$ over a wide range.
Benchmark.
At stationarity, the covariance $\Sigma(b)$ solves the Lyapunov equation
@@
so every stationary marginal variance scales as $1/b$. Importantly, $\eta$ does not appear in the equilibrium covariance; it governs mixing time, not the fixed point.
What the figure shows.
Figure 2 overlays simulated stationary variances with the Lyapunov prediction and shows clean $1/b$ scaling across the batch sweep. The flat coordinate mixes more slowly (as expected under weak curvature) but converges to the same equilibrium law. This serves as a baseline sanity check: the Fisher–Lyapunov equilibrium is not merely qualitative; it is quantitatively predictive. We observed early entry into the basin of Fisher geometry even from random starts; see Section 9.3 for a diagnostic based on the stability ratios $\delta_{F}$ and $\delta_{H}$.
@@
$d=2$, $H=\mathrm{diag}(1,0.1)$, $F=I_{2}$, 500 replicates per batch size, 200 000 steps each.
@@
Setup.
We keep the same sharp/flat curvature $H=\mathrm{diag}(1,0.1)$, but rotate an anisotropic noise-shape matrix relative to the curvature basis:
@@
with fixed batch size. This experiment isolates alignment as an operational parameter: the eigenvectors of $F$ move while $H$ is held fixed.
Benchmark.
Because the simulator is Euler–Maruyama with finite time step, the appropriate theoretical comparator is the discrete Lyapunov equation induced by the one-step dynamics. Writing $A=I-\eta H\Delta t$ and $Q=(2\eta/b)F(\theta)\Delta t$, the stationary covariance solves
@@
This removes discretization ambiguity and provides a clean, implementable target.
What the figure shows.
Figure 3 shows a clear transfer of stationary variance across coordinates as $\theta$ varies: when the noise eigenbasis rotates away from the curvature eigenbasis, equilibrium variance is redistributed in a predictable way. The discrete Lyapunov benchmark tracks the simulation closely across the sweep. The takeaway is structural: alignment governs which directions receive noise energy at equilibrium, even when scalar temperature $\eta/b$ is held fixed.
@@
$d=2$, $H=\mathrm{diag}(1,0.1)$, anisotropic $F(\theta)$ rotated over $0,\pi/2]$, 500 replicates per angle.
@@
Experiments 1–2 used $d=2$. We now move to $d=10$ to validate the Lyapunov equilibrium in a setting where anisotropy is richer. The curvature is $H=\mathrm{diag}(\lambda_{1},\ldots,\lambda_{10})$ with eigenvalues log-spaced between 1 and 100. The noise shape $G^{\star}$ is a random SPD matrix with eigenvalues log-spaced between 0.5 and 5; we also compare an isotropic control $G_{\mathrm{iso}}=\sigma^{2}I$ with $\sigma^{2}=\mathrm{Tr}(G^{\star})/d$ (trace-matched).
Setup.
Under constant step size $\eta=0.5/\lambda_{\max}(H)$ and batch size $b=64$, we run R=2, [^0] R=2{,}000 replicates and record the Fisher-metric risk $\mathbb{E}[\|\theta_{T}\|_{F^{\star}}^{2}]$ over increasing horizons $T\in\{300,800,2000,5000,10000\}$.
What the table shows.
Table 3 confirms the Lyapunov prediction (Corollary 5.8): both the anisotropic and isotropic empirical risks converge to the predicted stationary level as $T$ grows. The convergence ratio $\widehat{R}/R_{\infty}$ approaches unity from both sides, confirming that the discrete Lyapunov solution is the correct benchmark. The two Lyapunov plateaux nearly coincide because trace-matching is exact in this diagonal-$H$ setting; directional differences emerge only in higher-order structure (Claim 3 below).
@@
So what.
The Lyapunov equilibrium law gives the OR practitioner an exact closed-form prediction of steady-state risk from the problem primitives $(H,G^{\star},b)$ alone. Under a fixed compute budget, doubling $b$ halves the equilibrium covariance—but also halves the number of updates. The law makes this tradeoff precise.
@@ 8.2 Claim 2: Decaying Step Sizes Recover a $1/N$ Rate with a Geometry-Determined Constant
The second claim is that under a Robbins–Monro step-size schedule the Fisher-metric risk decays as $C/N$ where $N=Tb$ is the total oracle budget and the constant $C=\mathrm{Tr}(G^{\star}H^{-1})$ is pinned down by the intrinsic geometry. This validates the matching upper and lower bounds (Theorems 6.4 and 6.8).
@@
We switch to a decaying step size $\eta_{t}=\eta_{0}/(t+t_{0})$ with $\eta_{0}=1.2$, $t_{0}=50$, and the same $(H,G^{\star},G_{\mathrm{iso}})$ from Experiment 4. The oracle budget is $N=Tb$ with $b=64$.
What the figure shows.
Figure 4 (left) plots the Fisher-metric risk on log–log axes. Both anisotropic and isotropic risks decay as $1/N$ (fitted slopes $\approx-0.98$), confirming the upper bound of Theorem 6.4. The theoretical overlays $C/N$ with $C=\mathrm{Tr}(G^{\star}H^{-1})$ track the empirical curves closely. Figure 4 (right) shows the scaled risk $N\cdot\widehat{\mathrm{Risk}}$: the anisotropic curve stabilizes near $C_{\mathrm{aniso}}\approx 5.41$, the isotropic near $C_{\mathrm{iso}}\approx 5.07$, confirming the constant’s geometric origin. The inset marks the two theoretical constants $C=\mathrm{Tr}(G\,H^{-1})$.
@@
$\eta_{t}=1.2/(t+50)$, $b=64$, $R=800$ replicates per budget level.
Upper vs. lower bound comparison.
Table 4 summarizes the tightness of the bounds. The empirical risk is bounded between the information-theoretic lower bound (Theorem 6.8) and the upper bound (Theorem 6.4), with the scaled product $N\cdot\text{Risk}$ stabilizing close to $\mathrm{Tr}(G^{\star}H^{-1})$. At the largest budget (N=614,400N=614{,}400) the ratio of upper to lower bound constants is approximately $1.6\times$ for the anisotropic model, indicating near-tightness.
@@
The theoretical constant $C=\mathrm{Tr}(G^{\star}H^{-1})\approx 5.41$ (anisotropic) and $C\approx 5.07$ (isotropic); $N\cdot\mathrm{Risk}$ stabilizes above $d=10$ as expected.
So what.
The rate and constant are both determined by $(G^{\star},H^{\star})$; no tuning parameters remain once the geometry is identified (Corollary 6.10). For an OR practitioner, this means the oracle budget needed to achieve a target risk level can be computed in closed form from the problem’s Fisher/Godambe structure.
@@ 8.3 Claim 3: Scalar Temperature Matching Fails—Directionality Is the Signature of Geometry
The third claim is that matching the total noise power (scalar temperature $\tau=\eta/b$) is insufficient: the directional distribution of noise energy is a first-class observable that distinguishes the Fisher/Godambe-geometric model from any isotropic surrogate.
Why scalar temperature matching fails.
Experiments 4–5 show that the total Fisher-metric risk is well predicted by both the anisotropic and the trace-matched isotropic models. This is expected: the Lyapunov trace identity $\mathrm{Tr}(H^{\star}\Sigma_{\infty})=\frac{\tau}{2}\mathrm{Tr}(G^{\star})$ depends only on $\mathrm{Tr}(G^{\star})$, which is the same under trace matching. The isotropic surrogate gets the total right by construction. But the distribution of risk across directions is entirely wrong: the isotropic model spreads residual error uniformly, whereas the anisotropic model concentrates it along the top noise eigendirections. This directional mismatch is the operational signature of Fisher/Godambe geometry.
8.3.1 Experiment 3: Geometry vs. Scalar Temperature (“Rotation” Smoking Gun).
Setup.
A common practice in diffusion-based explanations of SGD is to match only a scalar noise level (“temperature”) while ignoring anisotropy. To show why this fails, we compare:
$\quad d\theta_{t}=-\eta H\theta_{t}\,dt+\sqrt{2\eta/b}\,F_{\star}(\phi)^{1/2}\,dW_{t},$
(Isotropic trace-matched)
$\quad d\theta_{t}=-\eta H\theta_{t}\,dt+\sqrt{2\eta/b}\,\sigma(\phi)\,dW_{t},\quad\sigma(\phi)^{2}:=\tfrac{1}{d}\mathrm{Tr}\!\bigl(F_{\star}(\phi)\bigr),$
where $F_{\star}(\phi)$ is anisotropic and rotates with $\phi$. Both models inject the same total noise power, but only the geometric model preserves directional structure.
Benchmark and diagnostic.
At stationarity, the geometric model has covariance $\Sigma_{F}(\phi)$ satisfying $H\Sigma_{F}+\Sigma_{F}H=(2/b)F_{\star}(\phi)$, while the isotropic surrogate has $H\Sigma_{I}+\Sigma_{I}H=(2/b)\sigma(\phi)^{2}I$. The diagnostic implication is immediate: $\Sigma_{I}(\phi)$ remains axis-aligned in the curvature basis, whereas $\Sigma_{F}(\phi)$ generically rotates (producing nontrivial cross-covariance and tilted level sets).
What the figures show.
Three diagnostics confirm the structural mismatch. Figure 5 visualizes the contrast directly via stationary covariance ellipses at three angles: the Fisher-geometric equilibrium tilts with $\varphi$, while the trace-matched isotropic surrogate remains axis-aligned. Figure 6 quantifies this by plotting the cross-covariance $\Sigma_{12}$ against $\varphi$: the Fisher model produces nontrivial off-diagonal entries that track theory, while the isotropic model stays identically zero. This is a sharp diagnostic—for any $\varphi\neq 0$, the geometric model produces $(\Sigma_{F})_{12}\neq 0$ while $(\Sigma_{I})_{12}=0$. Figure 7 shows how the marginal variances $\Sigma_{11}$ and $\Sigma_{22}$ redistribute across coordinates as $\varphi$ increases, again matching the Lyapunov prediction. The conclusion is crisp: temperature matching alone cannot reproduce Fisher/Godambe-structured equilibria; the full covariance geometry matters. This geometric discrepancy cannot be corrected by scalar temperature matching, since the off-diagonal structure encodes directional risk allocation that isotropic surrogates structurally lack. In information-theoretic terms, the two noise models are not equivalent experiments: the isotropic surrogate reproduces only the scalar temperature (trace) and cannot generate the cross-covariance $\Sigma_{12}(\varphi)$ observed under Fisher/Godambe noise. Hence the surrogates differ on any linear functional aligned with the rotated eigendirections—no scalar reparametrisation bridges the gap.
Figure 5: Stationary covariance ellipses at three rotation angles: the Fisher/Godambe equilibrium tilts with $\varphi$; the isotropic trace-matched surrogate remains axis-aligned.
$d=2$, $H=\mathrm{diag}(1,0.1)$, $F_{\star}(\varphi)$ with eigenvalues $(1.5,0.5)$, $b=64$. Ellipses show 95% level sets of the Lyapunov solution.
@@
$d=2$, Euler–Maruyama, 6 seeds per angle, T=2,500T=2{,}500, burn-in=800{}=800.
@@
We now move to $d=10$ and ask whether anisotropy affects how risk is distributed across directions. For $k=1,2,3$ we project $\theta_{T}$ onto the top eigenvectors $v_{k}$ of $G^{\star}$ and compute the directional second moment $\mathbb{E}[\langle v_{k},\theta_{T}\rangle^{2}]$.
What the figure shows.
Figure 8 (left) displays directional variances on log–log axes: both models decay as $1/N$, but the anisotropic curve is systematically above the isotropic one along the top eigendirections. Figure 8 (right) plots the ratio $\mathbb{E}_{\mathrm{aniso}}/\mathbb{E}_{\mathrm{iso}}$. At the largest budget the ratios are $1.65$, $1.45$, and $1.26$ for $k=1,2,3$, mirroring the eigenvalue ordering $\lambda_{1}>\lambda_{2}>\lambda_{3}$ of $G^{\star}$. This confirms the paper’s central message: anisotropic noise geometry concentrates residual error along the top noise eigendirections, an effect that isotropic surrogates cannot capture.
@@
Same parameters as Experiment 5; projections onto the top $k=1,2,3$ eigenvectors of $G^{\star}$.
So what.
For OR applications (e.g., revenue management, inventory optimization), different parameter directions may map to economically distinct decisions. The directional amplification ratios ($1.65\times$, $1.45\times$, $1.26\times$) show that the Fisher/Godambe geometry identifies which decisions carry the most residual risk—information that a scalar temperature model discards.

### 03 Apr 2026 v2

@@ 8 Numerical Validation
We validate the theory through six diagnostic experiments summarized in Table 3. The experiments probe downstream consequences of the identified covariance $\mathrm{Cov}(g_{B})\approx b^{-1}G^{\star}$ (Theorem 4.3): if the identification is correct, the Lyapunov equilibrium, convergence rates, and directional variance structure must all agree with the predictions derived from $G^{\star}$. Appendix B additionally validates the covariance identity itself under misspecification and approximate exchangeability.
@@
3 (Rotation diagnostic) 2 $\phi\in[0,\pi/2]$ (theory)
@@ 8.1 Claim 1: The Lyapunov Equilibrium Law Is Quantitatively Predictive
Throughout, when the simulator uses finite-step (Euler–Maruyama or direct SGD) dynamics, the mathematically correct benchmark is the discrete Lyapunov equation (Proposition 5.9); the continuous-time OU prediction is its small-$\eta$ asymptotic reduction. The regime of approximation accuracy is governed by the product $\eta\lambda_{\max}(H^{\star})$: when this quantity is small (say $\leq 0.1$), the continuous and discrete benchmarks are nearly indistinguishable; for moderate stepsizes the discrete equation is the appropriate comparator.
@@
with a two-dimensional sharp/flat curvature $H=\mathrm{diag}(1,0.1)$ and isotropic noise $F=I_{2}$, sweeping $b$ over a wide range. At stationarity, the covariance $\Sigma(b)$ solves the Lyapunov equation
@@
so every stationary marginal variance scales as $1/b$; $\eta$ governs mixing time but not the fixed point. For the Euler–Maruyama simulation used here, the discrete Lyapunov equation (49) is the exact benchmark; the continuous equation (48) coincides at leading order because the stepsize satisfies $\eta\lambda_{\max}(H)\leq 0.05$.
Figure 2 overlays simulated stationary variances with the Lyapunov prediction and shows clean $1/b$ scaling across the batch sweep. The flat coordinate mixes more slowly (as expected under weak curvature) but converges to the same equilibrium law. This serves as a baseline check: the Fisher–Lyapunov equilibrium is quantitatively predictive, not merely qualitative. We observed early entry into the basin of Fisher geometry even from random starts; see Section 9.3 for a diagnostic based on the stability ratios $\delta_{F}$ and $\delta_{H}$.
@@
$d=2$, $H=\mathrm{diag}(1,0.1)$, $F=I_{2}$, 500 replicates, 200k steps.
@@
We keep the same curvature $H=\mathrm{diag}(1,0.1)$ but rotate an anisotropic noise-shape matrix relative to the curvature basis:
@@
with fixed batch size, isolating alignment as an operational parameter. Because the simulator is Euler–Maruyama with finite time step, the appropriate comparator is the discrete Lyapunov equation: writing $A=I-\eta H\Delta t$ and $Q=(2\eta/b)F(\theta)\Delta t$, the stationary covariance solves
@@
Figure 3 shows a clear transfer of stationary variance across coordinates as $\varphi$ varies: when the noise eigenbasis rotates away from the curvature eigenbasis, equilibrium variance is redistributed in a predictable way, and the discrete Lyapunov benchmark tracks the simulation closely. The takeaway is structural: alignment governs which directions receive noise energy at equilibrium, even when scalar temperature $\eta/b$ is held fixed.
@@
$d=2$, $H=\mathrm{diag}(1,0.1)$, 500 replicates per angle.
@@
Experiments 1–2 used $d=2$. We now move to $d=10$ to validate the Lyapunov equilibrium in a setting where anisotropy is richer. The curvature is $H=\mathrm{diag}(\lambda_{1},\ldots,\lambda_{10})$ with eigenvalues log-spaced between 1 and 100; the noise shape $G^{\star}$ is a random SPD matrix with eigenvalues in $[0.5,5]$, and we compare an isotropic control $G_{\mathrm{iso}}=\sigma^{2}I$ with $\sigma^{2}=\mathrm{Tr}(G^{\star})/d$ (trace-matched). Under constant step size $\eta=0.5/\lambda_{\max}(H)$ and batch size $b=64$, we run R=2, [^0] R=2{,}000 replicates and record the Fisher-metric risk $\mathbb{E}[\|\theta_{T}\|_{F^{\star}}^{2}]$ over increasing horizons $T\in\{300,800,2000,5000,10000\}$.
Table 4 confirms the Lyapunov prediction (Corollary 5.7): both the anisotropic and isotropic empirical risks converge to the predicted stationary level as $T$ grows, with convergence ratios $\widehat{R}/R_{\infty}$ approaching unity from both sides. The two plateaux nearly coincide because trace-matching is exact in this diagonal-$H$ setting; directional differences emerge only in higher-order structure (Claim 3 below).
@@ 8.2 Claim 2: Decaying Step Sizes Recover a $1/N$ Rate with a Geometry-Determined Constant
The second claim is that under a Robbins–Monro step-size schedule the Fisher-metric risk decays as $C/N$ where $N=Tb$ is the total oracle budget and the constant $C$ is determined by the intrinsic geometry $(G^{\star},H^{\star},F^{\star})$. Specifically, the SA central-limit theorem gives an asymptotic covariance $V$ satisfying $({\eta_{0}H^{\star}-\tfrac{1}{2}I})V+V({\eta_{0}H^{\star}-\tfrac{1}{2}I})=\frac{\eta_{0}^{2}}{b}\,G^{\star}$, and $N\cdot\mathrm{Risk}\to b\,\mathrm{Tr}(F^{\star}V)$. This validates the matching upper and lower bounds (Theorem 6.5 and Proposition 6.10).
@@
We switch to a decaying step size $\eta_{t}=\eta_{0}/(t+t_{0})$ with $\eta_{0}=1.2$, $t_{0}=50$, and the same $(H,G^{\star},G_{\mathrm{iso}})$ from Experiment 4, with oracle budget $N=Tb$ and $b=64$. Figure 4 (left) plots the Fisher-metric risk on log–log axes: both anisotropic and isotropic risks decay as $1/N$ (fitted slopes $\approx-0.98$), confirming Theorem 6.5. The right panel shows the scaled risk $N\cdot\widehat{\mathrm{Risk}}$, which stabilizes near $\approx 14.2$ (anisotropic) and $\approx 14.0$ (isotropic)—both close to the SA–CLT prediction $b\,\mathrm{Tr}(F^{\star}V)$ where $V$ solves the Lyapunov equation above. The inset marks the auxiliary constant $\mathrm{Tr}(G^{\star}H^{-1})\approx 5.41$, which governs the constant-step OU stationary risk (Corollary 5.7); the ratio of the decaying-step plateau to this constant reflects the step-size schedule and the interplay between curvature and noise geometry.
@@
$\eta_{t}=1.2/(t+50)$, $b=64$, 800 replicates.
Table 5 summarizes the tightness of the bounds. The empirical risk is bounded between the information-theoretic lower bound (Proposition 6.10) and the upper bound (Theorem 6.5), with $N\cdot\text{Risk}$ stabilizing near $14.2$ (anisotropic) and $14.0$ (isotropic), both above the van Trees lower bound cd=10c\,d=10. At the largest budget (N=614,400N=614{,}400) the ratio of empirical constant to lower bound is approximately $1.4\times$, indicating near-tightness. The rate and constant are thus determined by $(G^{\star},H^{\star})$, so the oracle budget for a target risk can be computed from the problem’s intrinsic geometry.
@@
The SA–CLT Lyapunov prediction gives $b\,\mathrm{Tr}(F^{\star}V)\approx 14.2$ (anisotropic) and $\approx 14.0$ (isotropic); $N\cdot\mathrm{Risk}$ stabilizes above the van Trees lower bound cd=10c\,d=10 as expected.
@@ 8.3 Claim 3: Scalar Temperature Matching Fails—Directionality Is the Signature of Geometry
The third claim is that matching the total noise power (scalar temperature $\tau=\eta/b$) is insufficient: the directional distribution of noise energy is a distinguishing signature of the identified geometric model that no isotropic surrogate can reproduce. The Lyapunov trace identity $\mathrm{Tr}(H^{\star}\Sigma_{\infty})=\frac{\tau}{2}\mathrm{Tr}(G^{\star})$ depends only on $\mathrm{Tr}(G^{\star})$, so a trace-matched isotropic surrogate reproduces total risk by construction; but the directional distribution is entirely wrong, spreading residual error uniformly rather than concentrating it along top noise eigendirections.
8.3.1 Experiment 3: Geometry vs. Scalar Temperature (Rotation Diagnostic).
We compare a geometric OU model (diffusion covariance $\propto F_{\star}(\phi)$, rotating with angle $\phi$) against a trace-matched isotropic surrogate ($\sigma^{2}I$ with $\sigma^{2}=\mathrm{Tr}(F_{\star})/d$). At stationarity, the geometric covariance $\Sigma_{F}(\phi)$ generically rotates (producing nontrivial cross-covariance), while the isotropic $\Sigma_{I}(\phi)$ remains axis-aligned. Figures 57 confirm this: the ellipses tilt, cross-covariance $\Sigma_{12}$ tracks theory, and marginal variances redistribute—effects that temperature matching alone cannot capture.
Figure 5: Stationary covariance ellipses at three rotation angles: the identified equilibrium tilts with $\varphi$; the isotropic trace-matched surrogate remains axis-aligned.
$d=2$, $b=64$. Ellipses: 95% level sets of the Lyapunov solution.
@@
$d=2$, 6 seeds, T=2,500T=2{,}500, burn-in=800{}=800.
@@
We now move to $d=10$ and ask whether anisotropy affects how risk is distributed across directions. For $k=1,2,3$ we project $\theta_{T}$ onto the top eigenvectors $v_{k}$ of $G^{\star}$ and compute the directional second moment $\mathbb{E}[\langle v_{k},\theta_{T}\rangle^{2}]$. Figure 8 (left) displays these on log–log axes: both models decay as $1/N$, but the anisotropic curve is systematically above the isotropic one along the top eigendirections. The right panel plots the ratio $\mathbb{E}_{\mathrm{aniso}}/\mathbb{E}_{\mathrm{iso}}$; at the largest budget the ratios are $1.65$, $1.45$, and $1.26$ for $k=1,2,3$, mirroring the eigenvalue ordering $\lambda_{1}>\lambda_{2}>\lambda_{3}$ of $G^{\star}$. These amplification ratios show that the identified geometry determines which parameter directions carry the most residual risk—information that scalar temperature models discard.
@@
Same as Experiment 5; top $k=1,2,3$ eigenvectors of $G^{\star}$.